OpenStudy (anonymous):

ƒ(x) = √(x²+9) Is this a semi circle? :S

OpenStudy (anonymous):

no it's not. you see y=√(x²+9)>0, also y^2/9-x^2/9=1, therefore the graph is half of the hyperbola whose focus is (0,3√2) and which is above the x-axis.

OpenStudy (anonymous):

have I made you understood??

OpenStudy (anonymous):

where you get y^2/9-x^2/9=1 from? is that equation for hyperbola?

OpenStudy (anonymous):

square y then move it over so that you have 9 = y^2-x^2 then divide by nine to get 1=y^2/9-x^2/9

OpenStudy (anonymous):

oh i see now, i was thinking the x was in the denominator for y² because of computer writing ;)

OpenStudy (anonymous):

haha no, ken is just showing you that the equation is actually a hyperbola, and he proved it by simplifying

OpenStudy (anonymous):

yes. it;s like that. a standard hyperbola equation is x^2/a^2-y^2/b^2=1 or y^2/a^2-x^2/b^2=1

OpenStudy (anonymous):

the former has its foci on the x-axis, and the latter has its foci on the y-axis.

OpenStudy (anonymous):

kk, all good now. cheers

OpenStudy (anonymous):

cheers(^^)

OpenStudy (anonymous):

www.aceyourcollegeclasses.com is a great place to post your questions